# Rate-Splitting Multiple Access for Secure Near-Field Integrated Sensing and Communication

### v1

@@ I-B Main Contributions
Notations: Boldface upper-case letters, boldface lower-case letters, and calligraphy letters denote matrices, vectors, and sets. The complex matrix space of $N\times K$ dimensions is denoted by $\mathbb{C}^{N\times K}$. Superscripts ${(\bullet)}^{T}$ and ${(\bullet)}^{H}$ represent the transpose and Hermitian transpose, respectively. $\text{Re}\left(\bullet\right)$, $\text{Tr}\left(\bullet\right)$, and $\mathbb{E}\left[\bullet\right]$ denote the real part, trace, and statistical expectation. The Frobenius norm of matrix $\mathbf{X}$ is denoted by $\mathbf{X}||_{F}$. $\odot$ and $\otimes$ denote the Hadamard and Kronecker products. Variable $x\sim\mathcal{CN}(\mu,\sigma^{2})$ is a circularly symmetric complex Gaussian (CSCG) with mean $\mu$ and variance $\sigma^{2}$.
@@ II-B1 Communication performance
where $n_{k}\sim\mathcal{CN}\left(0,\sigma^{2}_{k}\right)$ and $n_{\text{e}}\sim\mathcal{CN}\left(0,\sigma^{2}_{\text{e}}\right)$ represent additional white Gaussian noise (AWGN). To facilitate the interpretation of the achievable rates for both legitimate users and the eavesdropper, we next provide the average received power at the $k$-th communication user and the target, which are respectively given by
@@ III-A Optimization problem reformulation
The coupling between analog beamfocuser $\mathbf{F}$ and digital beamfocuser $\mathbf{W}$ in (21) makes direct optimization intractable. Notably, $\mathbf{F}$ and $\mathbf{W}$ always appear in product form, so we introduces a fully digital beamfocuser $\mathbf{P}=\left[\mathbf{p}_{0},\mathbf{p}_{1},\dots,\mathbf{p}_{K}\right]\in\mathbb{C}^{N\times(K+1)}$ as an auxiliary variable, where $\mathbf{p}_{k}=\mathbf{Fw}_{k}$ for $\forall k\in\mathcal{K}_{1}=\{0,1,\dots,K\}$. By substituting $\mathbf{p}_{k}=\mathbf{Fw}_{k}$ into (2), the average received power at the $k$-th communication user and the target can be reformulated as
@@
$\text{CRB}\left(\theta_{\text{e}}\right)=\frac{\sigma^{2}_{\text{e}}\text{Tr}\left(\tilde{\mathbf{G}}\mathbf{PP}^{H}\tilde{\mathbf{G}}^{H}\right)}{2T\left|\tilde{\beta}_{\text{e}}\right|^{2}\left(\text{Tr}\left(\dot{\mathbf{G}}_{\theta_{\text{e}}}\mathbf{PP}^{H}\dot{\mathbf{G}}^{H}_{\theta_{\text{e}}}\right)\text{Tr}\left(\tilde{\mathbf{G}}\mathbf{PP}^{H}\tilde{\mathbf{G}}^{H}\right)-\left|\text{Tr}\left(\tilde{\mathbf{G}}\mathbf{PP}^{H}\dot{\mathbf{G}}^{H}_{\theta_{\text{e}}}\right)\right|^{2}\right)},$ (23)
@@ III-E Overall algorithm
  • Convergence: With any feasible initialization, both the inner and outer loops of Algorithm 2 generate non-decreasing objective values. In particular, each update of $\mathbf{P}$, $\mathbf{F}$, and $\mathbf{W}$ either improves or maintains the secrecy rate under a fixed penalty factor $\rho$. Since the secrecy rate is naturally upper-bounded, the inner loop is guaranteed to converge within finite iterations. Moreover, the outer loop is based on a penalty method, which has been established to converge to a stationary point [9120361]. Hence, the proposed algorithm can be regarded as convergent.
@@ Appendix B Reformulation of the CRB Matrix
and (77) as shown at the top of this page.
$\mathbf{J}_{12}\mathbf{J}^{T}_{12}$ $=\left[\begin{array}[]{cc}\text{Re}\left(\tilde{\beta}^{*}_{\text{e}}\text{Tr}\left(\tilde{\mathbf{G}}\mathbf{PP}^{H}\dot{\mathbf{G}}^{H}_{\theta_{\text{e}}}\right)\right)&-\text{Im}\left(\tilde{\beta}^{*}_{\text{e}}\text{Tr}\left(\tilde{\mathbf{G}}\mathbf{PP}^{H}\dot{\mathbf{G}}^{H}_{\theta_{\text{e}}}\right)\right)\\ \text{Re}\left(\tilde{\beta}^{*}_{\text{e}}\text{Tr}\left(\tilde{\mathbf{G}}\mathbf{PP}^{H}\dot{\mathbf{G}}^{H}_{r_{\text{e}}}\right)\right)&-\text{Im}\left(\tilde{\beta}^{*}_{\text{e}}\text{Tr}\left(\tilde{\mathbf{G}}\mathbf{PP}^{H}\dot{\mathbf{G}}^{H}_{r_{\text{e}}}\right)\right)\\ \end{array}\right]\left[\begin{array}[]{cc}\text{Re}\left(\tilde{\beta}^{*}_{\text{e}}\text{Tr}\left(\tilde{\mathbf{G}}\mathbf{PP}^{H}\dot{\mathbf{G}}^{H}_{\theta_{\text{e}}}\right)\right)&\text{Re}\left(\tilde{\beta}^{*}_{\text{e}}\text{Tr}\left(\tilde{\mathbf{G}}\mathbf{PP}^{H}\dot{\mathbf{G}}^{H}_{r_{\text{e}}}\right)\right)\\ -\text{Im}\left(\tilde{\beta}^{*}_{\text{e}}\text{Tr}\left(\tilde{\mathbf{G}}\mathbf{PP}^{H}\dot{\mathbf{G}}^{H}_{\theta_{\text{e}}}\right)\right)&-\text{Im}\left(\tilde{\beta}^{*}_{\text{e}}\text{Tr}\left(\tilde{\mathbf{G}}\mathbf{PP}^{H}\dot{\mathbf{G}}^{H}_{r_{\text{e}}}\right)\right)\\ \end{array}\right]$ (77)
$=\left|\tilde{\beta}_{\text{e}}\right|^{2}\left[\begin{array}[]{cc}\left|\text{Tr}\left(\tilde{\mathbf{G}}\mathbf{PP}^{H}\dot{\mathbf{G}}^{H}_{\theta_{\text{e}}}\right)\right|^{2}&\text{Re}\left(\text{Tr}\left(\tilde{\mathbf{G}}\mathbf{PP}^{H}\dot{\mathbf{G}}^{H}_{\theta_{\text{e}}}\right)\text{Tr}\left(\tilde{\mathbf{G}}\mathbf{PP}^{H}\dot{\mathbf{G}}^{H}_{r_{\text{e}}}\right)^{*}\right)\\ \text{Re}\left(\text{Tr}\left(\tilde{\mathbf{G}}\mathbf{PP}^{H}\dot{\mathbf{G}}^{H}_{\theta_{\text{e}}}\right)\text{Tr}\left(\tilde{\mathbf{G}}\mathbf{PP}^{H}\dot{\mathbf{G}}^{H}_{r_{\text{e}}}\right)^{*}\right)&\left|\text{Tr}\left(\tilde{\mathbf{G}}\mathbf{PP}^{H}\dot{\mathbf{G}}^{H}_{r_{\text{e}}}\right)\right|^{2}\\ \end{array}\right]$ (80)
@@
Substituting (81) and (77) into (20), we can obtain equation (82). When the target distance $r_{\text{e}}$ is known, $\dot{\mathbf{G}}_{r_{\text{e}}}=0$, and thus $\text{CRB}\left(r_{\text{e}},\theta_{\text{e}}\right)$ reduces to a matrix with only the upper-left element being nonzero, from which (23) holds directly. Conversely, when the target angle is known, $\dot{\mathbf{G}}_{\theta_{\text{e}}}=0$, and $\text{CRB}\left(r_{\text{e}},\theta_{\text{e}}\right)$ reduces to a matrix with only the lower-right element being nonzero, yielding (24).
$\text{CRB}\left(r_{\text{e}},\theta_{\text{e}}\right)=\frac{\sigma^{2}_{\text{e}}\text{Tr}\!\left(\tilde{\mathbf{G}}\mathbf{PP}^{H}\tilde{\mathbf{G}}^{H}\right)}{2T\left|\tilde{\beta}_{\text{e}}\right|^{2}}\left[\begin{array}[]{c|c}\begin{aligned} &\text{Tr}\!\left(\dot{\mathbf{G}}_{\theta_{\text{e}}}\mathbf{PP}^{H}\dot{\mathbf{G}}^{H}_{\theta_{\text{e}}}\right)\text{Tr}\!\left(\tilde{\mathbf{G}}\mathbf{PP}^{H}\tilde{\mathbf{G}}^{H}\right)\\ &\quad-\left|\text{Tr}\!\left(\tilde{\mathbf{G}}\mathbf{PP}^{H}\dot{\mathbf{G}}^{H}_{\theta_{\text{e}}}\right)\right|^{2}\end{aligned}&\begin{aligned} &\text{Re}\!\left(\text{Tr}\!\left(\dot{\mathbf{G}}_{r_{\text{e}}}\mathbf{PP}^{H}\dot{\mathbf{G}}^{H}_{\theta_{\text{e}}}\right)\text{Tr}\!\left(\tilde{\mathbf{G}}\mathbf{PP}^{H}\tilde{\mathbf{G}}^{H}\right)\right.\\ &\quad\left.-\text{Tr}\!\left(\tilde{\mathbf{G}}\mathbf{PP}^{H}\dot{\mathbf{G}}^{H}_{\theta_{\text{e}}}\right)\text{Tr}\!\left(\tilde{\mathbf{G}}\mathbf{PP}^{H}\dot{\mathbf{G}}^{H}_{r_{\text{e}}}\right)^{*}\right)\end{aligned}\\ \hline\cr\begin{aligned} &\text{Re}\!\left(\text{Tr}\!\left(\dot{\mathbf{G}}_{\theta_{\text{e}}}\mathbf{PP}^{H}\dot{\mathbf{G}}^{H}_{r_{\text{e}}}\right)\text{Tr}\!\left(\tilde{\mathbf{G}}\mathbf{PP}^{H}\tilde{\mathbf{G}}^{H}\right)\right.\\ &\quad\left.-\text{Tr}\!\left(\tilde{\mathbf{G}}\mathbf{PP}^{H}\dot{\mathbf{G}}^{H}_{\theta_{\text{e}}}\right)\text{Tr}\!\left(\tilde{\mathbf{G}}\mathbf{PP}^{H}\dot{\mathbf{G}}^{H}_{r_{\text{e}}}\right)^{*}\right)\end{aligned}&\begin{aligned} &\text{Tr}\!\left(\dot{\mathbf{G}}_{r_{\text{e}}}\mathbf{PP}^{H}\dot{\mathbf{G}}^{H}_{r_{\text{e}}}\right)\text{Tr}\!\left(\tilde{\mathbf{G}}\mathbf{PP}^{H}\tilde{\mathbf{G}}^{H}\right)\\ &\quad-\left|\text{Tr}\!\left(\tilde{\mathbf{G}}\mathbf{PP}^{H}\dot{\mathbf{G}}^{H}_{r_{\text{e}}}\right)\right|^{2}\end{aligned}\end{array}\right]^{\dagger}$ (82)

### v2

@@ I-B Main Contributions
Notations: Boldface upper-case letters, boldface lower-case letters, and calligraphy letters denote matrices, vectors, and sets. The complex matrix space of $N\times K$ dimensions is denoted by $\mathbb{C}^{N\times K}$. Superscripts ${(\bullet)}^{T}$, ${(\bullet)}^{H}$, and $(\bullet)$$^{\dagger}$ represent the transpose, Hermitian transpose, and pseudo-inverse, respectively. $\text{Re}\left(\bullet\right)$, $\text{Tr}\left(\bullet\right)$, and $\mathbb{E}\left[\bullet\right]$ denote the real part, trace, and statistical expectation. The Frobenius norm of matrix $\mathbf{X}$ is denoted by $\mathbf{X}||_{F}$. $\odot$ and $\otimes$ denote the Hadamard and Kronecker products. Variable $x\sim\mathcal{CN}(\mu,\sigma^{2})$ is a circularly symmetric complex Gaussian (CSCG) with mean $\mu$ and variance $\sigma^{2}$.
@@ II-B1 Communication performance
where $n_{k}\sim\mathcal{CN}\left(0,\sigma^{2}_{k}\right)$ and $n_{\text{e}}\sim\mathcal{CN}\left(0,\sigma^{2}_{\text{e}}\right)$ represent additional white Gaussian noise (AWGN). To facilitate the interpretation of the achievable rates for both legitimate users and the eavesdropper, we next provide the average received power at the $k$-th communication user and the target, which are respectively given by [7555358]
@@ III-A Optimization problem reformulation
The coupling between analog beamfocuser $\mathbf{F}$ and digital beamfocuser $\mathbf{W}$ in (21) makes direct optimization intractable. Notably, $\mathbf{F}$ and $\mathbf{W}$ always appear in product form, so we introduces a fully digital beamfocuser $\mathbf{P}=\left[\mathbf{p}_{0},\mathbf{p}_{1},\dots,\mathbf{p}_{K}\right]\in\mathbb{C}^{N\times(K+1)}$ as an auxiliary variable, where $\mathbf{p}_{k}=\mathbf{Fw}_{k}$ for $\forall k\in\mathcal{K}_{1}=\{0,1,\dots,K\}$. By substituting $\mathbf{p}_{k}=\mathbf{Fw}_{k}$ into (2), the average received power at the $k$-th communication user and the target can be reformulated as [7555358]
@@
$\text{CRB}\left(\theta_{\text{e}}\right)=\frac{\sigma^{2}_{\text{e}}\text{Tr}\!\left(\tilde{\mathbf{G}}\mathbf{PP}^{H}\tilde{\mathbf{G}}^{H}\right)}{2T\left|\tilde{\beta}_{\text{e}}\right|^{2}\!\left(\text{Tr}\!\left(\dot{\mathbf{G}}_{\theta_{\text{e}}}\mathbf{PP}^{H}\dot{\mathbf{G}}^{H}_{\theta_{\text{e}}}\right)\text{Tr}\!\left(\tilde{\mathbf{G}}\mathbf{PP}^{H}\tilde{\mathbf{G}}^{H}\right)-\left|\text{Tr}\!\left(\tilde{\mathbf{G}}\mathbf{PP}^{H}\dot{\mathbf{G}}^{H}_{\theta_{\text{e}}}\right)\right|^{2}\right)}.$ (23)
@@ III-E Overall algorithm
  • Convergence: With any feasible initialization, both the inner and outer loops of Algorithm 2 generate non-decreasing objective values. In particular, each update of $\mathbf{P}$, $\mathbf{F}$, and $\mathbf{W}$ either improves or maintains the secrecy rate under a fixed penalty factor $\rho$. Since the secrecy rate is naturally upper-bounded, the inner loop is guaranteed to converge within finite iterations. Moreover, the outer loop is based on a penalty method, which has been established to converge to a stationary point [9120361]. Hence, the proposed penalty-based BCD algorithm can be regarded as convergent.
@@ Appendix B Reformulation of the CRB Matrix
and
$\mathbf{J}_{12}\mathbf{J}^{T}_{12}=$ $\left[\begin{array}[]{cc}\text{Re}\left(\tilde{\beta}^{*}_{\text{e}}\kappa_{1}\right)&-\text{Im}\left(\tilde{\beta}^{*}_{\text{e}}\kappa_{1}\right)\\ \text{Re}\left(\tilde{\beta}^{*}_{\text{e}}\kappa_{2}\right)&-\text{Im}\left(\tilde{\beta}^{*}_{\text{e}}\kappa_{2}\right)\\ \end{array}\right]\times$ (75)
$\left[\begin{array}[]{cc}\text{Re}\left(\tilde{\beta}^{*}_{\text{e}}\kappa_{1}\right)&\text{Re}\left(\tilde{\beta}^{*}_{\text{e}}\kappa_{2}\right)\\ -\text{Im}\left(\tilde{\beta}^{*}_{\text{e}}\kappa_{1}\right)&-\text{Im}\left(\tilde{\beta}^{*}_{\text{e}}\kappa_{2}\right)\\ \end{array}\right]$ (78)
$=$ $\left|\tilde{\beta}_{\text{e}}\right|^{2}\left[\begin{array}[]{cc}\left|\kappa_{1}\right|^{2}&\text{Re}\left(\kappa_{1}\kappa^{*}_{2}\right)\\ \text{Re}\left(\kappa_{1}\kappa^{*}_{2}\right)&\left|\kappa_{2}\right|^{2}\\ \end{array}\right]$ (81)
where the symbol $\times$ denotes matrix multiplication between two matrices, $\kappa_{1}=\text{Tr}\left(\tilde{\mathbf{G}}\mathbf{PP}^{H}\dot{\mathbf{G}}^{H}_{\theta_{\text{e}}}\right)$ and $\kappa_{2}=\text{Tr}\left(\tilde{\mathbf{G}}\mathbf{PP}^{H}\dot{\mathbf{G}}^{H}_{r_{\text{e}}}\right)$, respectively.
@@
Substituting (82) and (75) into (20), we can obtain
$\text{CRB}\left(r_{\text{e}},\theta_{\text{e}}\right)=\frac{\sigma^{2}_{\text{e}}\text{Tr}\!\left(\tilde{\mathbf{G}}\mathbf{PP}^{H}\tilde{\mathbf{G}}^{H}\right)}{2T\left|\tilde{\beta}_{\text{e}}\right|^{2}}\left[\begin{array}[]{c|c}C_{11}&C_{12}\\ \hline\cr C_{21}&C_{22}\\ \end{array}\right]^{\dagger},$ (85)
where
$C_{11}=$ $\text{Tr}\!\left(\dot{\mathbf{G}}_{\theta_{\text{e}}}\mathbf{PP}^{H}\dot{\mathbf{G}}^{H}_{\theta_{\text{e}}}\right)\text{Tr}\!\left(\tilde{\mathbf{G}}\mathbf{PP}^{H}\tilde{\mathbf{G}}^{H}\right)$
$-\left|\text{Tr}\!\left(\tilde{\mathbf{G}}\mathbf{PP}^{H}\dot{\mathbf{G}}^{H}_{\theta_{\text{e}}}\right)\right|^{2},$ (86a)
$C_{12}=$ $\text{Re}\!\left(\text{Tr}\!\left(\dot{\mathbf{G}}_{r_{\text{e}}}\mathbf{PP}^{H}\dot{\mathbf{G}}^{H}_{\theta_{\text{e}}}\right)\text{Tr}\!\left(\tilde{\mathbf{G}}\mathbf{PP}^{H}\tilde{\mathbf{G}}^{H}\right)\right$.
$\left.-\text{Tr}\!\left(\tilde{\mathbf{G}}\mathbf{PP}^{H}\dot{\mathbf{G}}^{H}_{\theta_{\text{e}}}\right)\text{Tr}\!\left(\tilde{\mathbf{G}}\mathbf{PP}^{H}\dot{\mathbf{G}}^{H}_{r_{\text{e}}}\right)^{*}\right),$ (86b)
$C_{21}=$ $\text{Re}\!\left(\text{Tr}\!\left(\dot{\mathbf{G}}_{\theta_{\text{e}}}\mathbf{PP}^{H}\dot{\mathbf{G}}^{H}_{r_{\text{e}}}\right)\text{Tr}\!\left(\tilde{\mathbf{G}}\mathbf{PP}^{H}\tilde{\mathbf{G}}^{H}\right)\right$.
$\left.-\text{Tr}\!\left(\tilde{\mathbf{G}}\mathbf{PP}^{H}\dot{\mathbf{G}}^{H}_{\theta_{\text{e}}}\right)\text{Tr}\!\left(\tilde{\mathbf{G}}\mathbf{PP}^{H}\dot{\mathbf{G}}^{H}_{r_{\text{e}}}\right)^{*}\right),$ (86c)
$C_{22}=$ $\text{Tr}\!\left(\dot{\mathbf{G}}_{r_{\text{e}}}\mathbf{PP}^{H}\dot{\mathbf{G}}^{H}_{r_{\text{e}}}\right)\text{Tr}\!\left(\tilde{\mathbf{G}}\mathbf{PP}^{H}\tilde{\mathbf{G}}^{H}\right)$
$-\left|\text{Tr}\!\left(\tilde{\mathbf{G}}\mathbf{PP}^{H}\dot{\mathbf{G}}^{H}_{r_{\text{e}}}\right)\right|^{2}.$ (86d)
When the target distance $r_{\text{e}}$ is known, $\dot{\mathbf{G}}_{r_{\text{e}}}=0$, and thus $\text{CRB}\left(r_{\text{e}},\theta_{\text{e}}\right)$ reduces to a matrix with only the upper-left element being nonzero, from which (23) holds directly. Conversely, when the target angle is known, $\dot{\mathbf{G}}_{\theta_{\text{e}}}=0$, and $\text{CRB}\left(r_{\text{e}},\theta_{\text{e}}\right)$ reduces to a matrix with only the lower-right element being nonzero, yielding (24).